OpenStudy (anonymous):

convergence of k(1/3)^k

OpenStudy (anonymous):

i actually just figured out

ganeshie8 (ganeshie8):

Okay great :)

OpenStudy (anonymous):

its geometric and bc r < 1 it converges right?

ganeshie8 (ganeshie8):

how do you know it is geometric with a \(k\) attached to it ?

ganeshie8 (ganeshie8):

\(\sum\limits_{k=0}^{\infty}(1/3)^k\) is easy to spot as geometric series but how do you know \(\sum\limits_{k=0}^{\infty} \color{Red}{k}(1/3)^k\) is related to geometric series ?

OpenStudy (anonymous):

i guess i don't?

ganeshie8 (ganeshie8):

It is easy to see the series converges by rewritign it like this : \[\sum\limits_{k=0}^{\infty} k(1/3)^k = \sum\limits_{k=0}^{\infty} \dfrac{k}{3^k}\] Notice the numerator is a polynomial and the denominator is an exponential function which grows wildly. You may use ratio test to conclude that the series converges

OpenStudy (anonymous):

yeah that makes a lot of sense i could use some more help ill post a different question in a second!

ganeshie8 (ganeshie8):

Sure :) there is a neat way to evaluate this sum if you're interested..

ganeshie8 (ganeshie8):

\[\sum\limits_{k=0}^{\infty} k(1/3)^{k}\] is same as\[\sum\limits_{k=0}^{\infty} nx^n\] where \(x = 1/3\)

ganeshie8 (ganeshie8):

and you know the geometric series \[\frac{1}{1-x} = \sum\limits_{k=0}^{\infty} x^n\] differentiate both sides and multiply by \(x\)

OpenStudy (anonymous):

i see why you're a 'honorary professor of mathematics'